OpenStudy (lochana):

Please someone help me with this "differential equation" problem. Prove that for any straight line through origin \(\frac{y}{x} = \frac{dy}{dx}\). and interpret this.

OpenStudy (lochana):

should I integrate it?

OpenStudy (dls):

Its simply saying that slope of a line = m = y/x = tan theta (if it passes through origin then c=0 in the standard equation of line y=mx+c)

OpenStudy (lochana):

I see.. but do you know how to prove it?

OpenStudy (dls):

\[\Large \frac{y}{x} = \frac{dy}{dx}\] \[\large \int\limits_0^y \frac{dy}{y} = \int\limits_0^x \frac{dx}{x}\]

OpenStudy (lochana):

yeah. then it becomes\(lnx = lny\).. now I can say it is y=x. right?

OpenStudy (dls):

You missed a step. Even though its obvious but let me state it. \[\Large \ln(y) = \ln(x) + c\] Since it passes through origin then (0,0) satisfies it. Hence c=0 Now you can simply take antilog on both the sides and get y=x :)

OpenStudy (lochana):

got it. @DLS Thanks for showing it. it really helped. bye

OpenStudy (baru):

Y=cx (where c is the constant of integration)

ganeshie8 (ganeshie8):

basically the differential equation \(dy/dx = y/x\) represents the family of straight lines passing thru origin

ganeshie8 (ganeshie8):

\(y = cx\tag{1}\) differentiating both sides with respect to x gives \(dy/dx =c\) from \((1)\) we have \(c = y/x\)

OpenStudy (baru):

or if you want to integrate, ln y =ln(x) +ln(c) ln(y)=ln(x*c) y=cx